# DesignBench: A Comprehensive Benchmark for MLLM-based Front-end Code Generation

### v2

@@ 2.1. Web Development Process
2.2. Front-end Web Development Framework
@@ 4.1. Task Definition
Design Generation ($\mathcal{T}_{G}$). The objective of design generation is to generate expected code based on the UI Mockups. Formally, given a UI design image $I$, the task aims to generate corresponding UI code $C$ such that $\mathcal{T}_{G}:I\rightarrow C$. The input contains the UI design image $I$, and the output is the UI code $C$ that accurately reproduces the visual layout and styling.
@@ 4.2. Data Collection
For websites without applying frameworks, we sample 120 webpages with different lengths from the webcode2m dataset (Gui et al., 2025). Ultimately, we compile a dataset consisting of 430 webpages. After getting the link of the webpage, we apply single-file tool (Lormeau, 2023) to save the webpage as a stand-alone file, then replace the images in the web page with placeholders, and finally use selenium (Selenium, 2025) to take screenshots for constructing the input design image $I$ for task $\mathcal{T}_{G}$.
@@ 5.4. Metric
Code Metrics. (1) Compilation Success Rate (CSR). This metric represents the percentage of generated code that compiles successfully without errors. Assume that the total number of samples is N and the number of samples compiled successfully is S, then $CSR=\frac{S}{N}$. (2) Code Modification Location Similarity (CMLS). We employ the Jaccard similarity (Thada and Jaglan, 2013) to quantify the precision of code modifications on design edit and design repair tasks by comparing the sets of code modification operations between the ground truth and generated code. We first parse the code into an Abstract Syntax Tree (AST), and then identify the operation (i.e., add, delete and modify) by the AST node comparison. Let $A$ represent the set of code modification operations in the ground truth code and $B$ represent the set of code modification operations in the generated code. The CMLS is formally defined as: $CMLS(A,B)=\frac{|A\cap B|}{|A\cup B|}$. (3) Code Modification Content Similarity (CMCS). While CMLS focuses on modification locations, CMCS evaluates the semantic correctness of modified code content. Given modification operation sets $A$ and $B$, we match operations based on AST node correspondence and denote the matched pairs as $M\subseteq A\times B$. For each $(a,b)\in M$, CodeBLEU (ren2020codebleu) is used to compute content similarity, while unmatched operations receive a score of zero. The CMCS is defined as: $CMCS(A,B)=\frac{1}{|A\cup B|}\sum_{(a,b)\in M}\text{CodeBLEU}(a,b)$.
@@ 6.3. RQ3: How does varying task difficulty impact MLLM performance?
For Design Generation task, which requires generating code from scratch, we calculate the complexity based on the image size $I$, UI elements number $U$, color variety $C$ (i.e., the number of colors in the UI) and layout complexity $L$ (i.e., number of layout methods + number of nesting levels + number of CSS layout attributes). We then normalize these scores to between 0 and 100. The complexity score: $S=0.25\times I+0.25\times U+0.25\times C+0.25\times L$. Scores greater than 80 are complex, less than 30 are simple, and between 30 and 80 are medium.
@@ 6.5.3. Issue Detection Limitation
6.6. RQ6: What mistakes do MLLMs make on the three tasks?

### v3

@@ 2.1. Web Development Process
$2.$ Front-end Web Development Framework
@@ 4.1. Task Definition
Design Generation ($\mathcal{T}_{G}$). The objective of design generation is to generate expected code based on the UI Mockups. Formally, given a UI design image II, the task aims to generate corresponding UI code $C$ such that $\mathcal{T}_{G}:I\rightarrow C$. The input contains the UI design image II, and the output is the UI code $C$ that accurately reproduces the visual layout and styling.
@@ 4.2. Data Collection
For websites without applying frameworks, we sample 120 webpages with different lengths from the webcode2m dataset (Gui et al., 2025). Ultimately, we compile a dataset consisting of 430 webpages. After getting the link of the webpage, we apply single-file tool (Lormeau, 2023) to save the webpage as a stand-alone file, then replace the images in the web page with placeholders, and finally use selenium (Selenium, 2025) to take screenshots for constructing the input design image II for task $\mathcal{T}_{G}$.
@@ 5.4. Metric
Code Metrics. (1) Compilation Success Rate (CSR). This metric represents the percentage of generated code that compiles successfully without errors. Assume that the total number of samples is N and the number of samples compiled successfully is S, then $CSR=\frac{S}{N}$. (2) Code Modification Location Similarity (CMLS). We employ the Jaccard similarity (Thada and Jaglan, 2013) to quantify the precision of code modifications on design edit and design repair tasks by comparing the sets of code modification operations between the ground truth and generated code. We first parse the code into an Abstract Syntax Tree (AST), and then identify the operation (i.e., add, delete and modify) by the AST node comparison. Let AA represent the set of code modification operations in the ground truth code and $B$ represent the set of code modification operations in the generated code. The CMLS is formally defined as: $CMLS(A,B)=\frac{|A\cap B|}{|A\cup B|}$. (3) Code Modification Content Similarity (CMCS). While CMLS focuses on modification locations, CMCS evaluates the semantic correctness of modified code content. Given modification operation sets AA and $B$, we match operations based on AST node correspondence and denote the matched pairs as $M\subseteq A\times B$. For each $(a,b)\in M$, CodeBLEU (ren2020codebleu) is used to compute content similarity, while unmatched operations receive a score of zero. The CMCS is defined as: $CMCS(A,B)=\frac{1}{|A\cup B|}\sum_{(a,b)\in M}\text{CodeBLEU}(a,b)$.
@@ 6.3. RQ3: How does varying task difficulty impact MLLM performance?
For Design Generation task, which requires generating code from scratch, we calculate the complexity based on the image size II, UI elements number $U$, color variety $C$ (i.e., the number of colors in the UI) and layout complexity $L$ (i.e., number of layout methods + number of nesting levels + number of CSS layout attributes). We then normalize these scores to between 0 and 100. The complexity score: $S=0.25\times I+0.25\times U+0.25\times C+0.25\times L$. Scores greater than 80 are complex, less than 30 are simple, and between 30 and 80 are medium.
@@ 6.5.3. Issue Detection Limitation
$6.$ RQ6: What mistakes do MLLMs make on the three tasks?